# Disorder to Order Transition in 1D Nonreciprocal Cahn-Hilliard Model

### v1

@@ Abstract
We extensively study the phenomenology of one dimensional Nonreciprocal Cahn Hilliard model for varying nonreciprocity $(\alpha)$ and different boundary conditions. At small $\alpha$, a perturbed uniform state evolves to defect laden configuration that lack global polar order. Defects are the sources and sinks of travelling waves and nonreciprocity selects defects with a unique wave number that increases monotonically with $\alpha$. A critical threshold $\alpha_{c}$ marks the onset of a transition to states with finite global polar order. For periodic boundaries, above $\alpha_{c}$, the system shows travelling waves that are completely ordered. In contrast, travelling waves are incompatible with Dirichlet and Neumann boundaries. Instead, for $\alpha\gtrsim\alpha_{c}$, we find fluctuating domains that show intermittent polar order and at large $\alpha$, the system partitions into two domains with opposite polar order.
@@ I Introduction
The Nonreciprocal Cahn-Hilliard (NRCH) model describes phase separation of multi-component mixtures in the presence of nonreciprocal couplings [1, 2, 3, 4, 5, 6, 7, 8]. Systems with nonreciprocity are intrinsically out of equilibrium and they show rich phenomenology [9, 10, 11, 12, 13, 14, 15, 16]. In the NRCH model, the spontaneous breaking of parity and time-reversal symmetries leads to the formation of travelling density bands [1, 2], suppressed coarsening dynamics [17, 1], localised states [18], chaotic steady states [8], true long-range polar order in two dimensions [7], and enhanced stability in multispecies mixtures [5]. Introduced phenomenologically, the NRCH model emerges as a universal amplitude equation for a conserved-Hopf instability in systems with two conservation laws [19] and can also be derived from a systematic coarse-graining of a microscopic model of phoretically active Janus colloids [20, 21].
@@
In Rana and Golestanian [3, 4], we presented defect dynamics of the NRCH model in two dimensions (2D). The model admits spirals with unit magnitude topological charge and topologically neutral targets. These defects are the sources of the travelling waves, and thus the precursors of global polar order. For a given strength of nonreciprocal coupling $\alpha$, defects with a unique asymptotic wave number, $k_{\infty}\propto\sqrt{\alpha}$ and amplitude $R_{\infty}=\sqrt{1-k_{\infty}^{2}}$, are selected [3]. At a critical threshold of nonreciprocity $\alpha_{c}$; we find a disorder to order transition. For $\alpha<\alpha_{c}$, random disordered states evolve to quasi-stationary defect networks with no global polar order. On the other hand, for $\alpha>\alpha_{c}$, we find travelling waves. Below $\alpha_{c}$, the topological composition of the defect networks is also $\alpha$ dependent. For $\alpha\ll\alpha_{c}$, we always find spirals, whereas for $\alpha\lesssim\alpha_{c}$ we primarily observe target networks. Our further study [4] on the interactions between a pair of defects revealed that the stability of targets fundamentally alters the nature of defect interactions and sets them apart from the previously studied out of equilibrium systems with a nonconserved order parameter [22, 23, 24, 25]. Thus, NRCH provides a fertile ground to investigate the defect phenomenology in conserved nonreciprocal systems.
In this paper, we present an extensive study on the defects of the NRCH model in one dimension (1D). Our main motivation is to investigate defect dynamics of the NRCH model in a simpler setting as compared to our previous 2D studies [3, 4]. Charged spirals are ruled out, and we observe stable sources and sinks, which are the direct analogue of two dimensional targets. These sources and sinks are arranged in an alternating configurations and their motion is limited to the line. Similar to 2D, sources and sinks with a particular wavenumber $k_{\infty}$ that increases with $\alpha$ are selected. The 1D NRCH model also shows a disorder-order transition. Crucially and in a stark contrast to 2D, this transition occurs at the crossover $\alpha$ predicted by the Eckhaus instability of the plane waves. We further investigate the features of the disorder-order transition for different boundary conditions. Since travelling waves are not accepted solutions of the Neumann or the Dirichlet boundary conditions, at $\alpha\gtrsim\alpha_{\times}$, the system shows intermittent behaviour, with fluctuating polar order for these boundary conditions. For larger $\alpha$, the system splits into subdomains with opposing polar order.
@@
The rest of the paper is organized as follows. In Section II, we present our model and methods. In Section III, we present our results for periodic boundary conditions. We highlight the features of the source and sink solutions, wave number selection, defect density, polar order, and the disorder to order transition. In Section IV, we consider Dirichlet and Neumann boundary conditions and discuss how the boundary conditions alter the phenomenology. We conclude the paper with a discussion and future perspectives.
@@ II Model and Methods
We consider a minimal model of two conserved scalar fields $\phi_{1}(x,t)$ and $\phi_{2}(x,t)$ with nonreciprocal interactions on a one dimensional domain. The conserved dynamics of the complex scalar order parameter $\phi=\phi_{1}+i\phi_{2}$ is governed by the continuity equation $\partial_{t}\phi=\partial_{x}^{2}\mu$, where $\mu=\delta\mathcal{F}/\delta\phi^{*}+i\alpha\phi$ is the non-equilibrium chemical potential. The equilibrium contribution to $\mu$ promotes phase-separation and is derived from the free energy functional $\mathcal{F}=\int\mathop{}\!\mathrm{d}x\left(-|\phi|^{2}/2+|\phi|^{4}/4+|\nabla\phi|^{2}/2\right)$. The nonequilibrium part encodes the nonreciprocal interactions between the two species; quantified by the parameter $\alpha$. For $\alpha=0$, the model describes equilibrium phase-separation of two interacting species, and an initial disordered state coarsens towards a bulk-phase separated state. Any $\alpha\neq 0$ sets the system out of equilibrium and it describes nonreciprocal phase-separation. A positive $\alpha$ implies that $\phi_{1}$ chases $\phi_{2}$, and the model is invariant under the simultaneous transformation $\alpha\to-\alpha$ and $\phi_{1,2}\to\phi_{2,1}$. Combining everything together, we obtain the following non-dimensional equation for the evolution of $\phi(x,t)$ [3, 4, 1, 2]
@@
which reveals that $\forall~|k|<1$ the perturbations grow in an oscillatory manner. The behaviour of the system depends on $\alpha$, domain length $L$, and the choice of boundary conditions. In this manuscript, we systematically study the phase-behaviour of Eq. 1 with varying strength of nonreciprocity $(\alpha)$ for various domain sizes $(L)$ and boundary conditions. We consider three different kinds of boundary conditions that are relevant to the phenomenon of phase separation [26]:
@@
    No-flux boundary conditions (N-BC), where
    $\partial_{x}\phi|_{0}=\partial_{x}\phi|_{L}=0~\mathrm{and}~\partial_{x}\mu|_{0}=\partial_{x}\mu|_{L}=0.$
@@
    Dirichlet boundary conditions (D-BC), where
    $\phi(0)=\phi(L)=0~\mathrm{and}~\mu(0)=\mu(L)=0.$
To study the phenomenology, we numerically integrate Eq. 1 on a 1D domain of length $L$ that is discretized using $N$ points. For P-BC, we use a pseudo-spectral algorithm that employs a second-order exponential time-differencing scheme [27] for time marching. For N-BC and D-BC, we perform the simulations using a Chebyshev polynomials based spectral Tau method implemented using the Dedalus framework [28]. Additionally, we also use Dedalus to cross-check our results for the periodic domain. The source code to simulate the 1D NRCH model using Dedalus is available at https://github.com/navdeeprana/DedalusScripts.
We start all our simulations with a small perturbation to the homogeneous state. i.e., $\phi(x,0)=\delta\phi(x)$, where at every point $x$ the initial perturbation $\delta\phi(x)$ is drawn from a uniform distribution. Further, we impose $\int_{0}^{L}\phi(x,0)\mathop{}\!\mathrm{d}x=0$, which remains constant throughout the entire evolution due to the overall conservation law. Unless reported otherwise, we average over eight or more independent realizations of initial conditions for each data point reported in the paper.
Figure 1: Kymograph plot of the polar order parameter $J(x,t)$ highlighting the evolution of disordered states. Defects are located on zeroes of $J(x,t)$, which appear white in the kymograph. We also mark the exact defect positions (small black points) at late times. Initially, numerous source-sink pairs are formed, which merge and the system settles into a stable configuration, which persists till the end of the simulations.
Figure 2: Sources and sinks in the 1D NRCH model. (a, b) Representative plot of $\real{\phi}$ and the polar order parameter $J$ (normalised by it’s maximum value), respectively. Sources (blue circles) and sinks (orange circles) are arranged in an alternating manner on the 1D domain. $J$ vanishes at defect centres and is constant far from defects, where the wavefront is that of a plain wave. (c, d) Plot of $R(x)$ and $k(x)$ for sources (left column) and sinks (right column) for different values of $\alpha$. For these plots, $x$ is measured from the centre of the defects. At defect centres, $\partial_{x}R$ and $k(x)$ vanish. Far from the defects, $R(x)$ and $k(x)$ approach their plain wave forms.
@@ III Periodic boundary conditions
For P-BC, the NRCH model Eq. 1 admits travelling wave solutions of the form
@@
where $k<1$, $R^{2}=1-k^{2}$, and $\omega=\alpha k^{2}$. For small $k$, travelling states are stable to small perturbations, however an Eckhaus instability kicks in at large $k$ and restricts the allowed wave number range to $0\leq k^{2}<1/3$ [23, 8, 3, 29]. In our previous studies [3, 4], we have shown that in 2D, at low $\alpha$, disordered states evolves to defect configurations composed of spirals and targets. In 1D, we find similar phenomenology. A random state evolves into multi-defect configurations that persist up to the end of simulation time (see Fig. 1 and Fig. 2). In what follows, we discuss the properties of these states in detail.
@@ III.1 Sources and Sinks
$\phi(x,t)=R(x)\,e^{i\left[Z(x)-\omega t\right]},$ (4)
@@
where $x$ is measured from defect location, $R(x)$ is the amplitude, $Z(x)$ is the phase, and $k(x)\equiv\mathop{}\!\mathrm{d}Z/\mathop{}\!\mathrm{d}x$. They are the sources and sinks of travelling waves and we shall collectively call them defects. In Fig. 2(a), we plot a representative multi-defect configuration obtained from our P-BC simulations. The sources and sinks are arranged in an alternating manner on the 1D domain. Furthermore, for a periodic domain, the total number of sources is always equal to the total number of sinks. It is easy to identify the position of these defects by using the polar order parameter
$J(x,t)=\mathrm{Im}(\phi^{*}\partial_{x}\phi)=\frac{1}{2i}\left(\phi^{*}\partial_{x}\phi-\phi\partial_{x}\phi^{*}\right).$ (5)
@@
As shown in Fig. 2(b), at a source (sink), $J$ vanishes with a positive (negative) slope, thus at a given time $t$, the number of defects is equal to the number of zeroes of $J(x,t)$. Substituting the defect ansatz in the definition of $J(x,t)$, we further find that at the defect core $\partial_{x}(R^{2})=\partial_{x}(\phi\bar{\phi})=2\bar{\phi}\partial_{x}\phi=2\phi\partial_{x}\bar{\phi}$. Thus, at the defect positions,
$\partial_{x}R(x)=0,~\mathrm{and}~k(x)=0.$ (6)
@@
In Fig. 2(c,d), we plot the profiles of $R(x)$ and $k(x)$ for sources and sinks for different $\alpha$, where $x$ is measured from the location of the defect. For a source, $k(x)>0(<0)$ when $x>0(<0)$, and vice versa for a sink. It is evident that R(x)=R(x)R(-x)=R(x) and k(x)=k(x)k(-x)=-k(x), which implies that for small $x$, $R(x)\sim a_{0}-a_{2}x^{2}$ and $k(x)\sim b_{1}x-b_{3}x^{3}$. Far away from the defect centre, the wavefront approaches that of a travelling wave (see Fig. 2(a)). Thus, when $x\to\pm\infty$, $k(x)\to\pm k_{\infty}$ for a source and $k(x)\to\mp k_{\infty}$ for a sink. The amplitude $R(x)\to\sqrt{1-k_{\infty}^{2}}$ also becomes constant for both and to ensure proper oscillations, we require $\omega=\alpha k_{\infty}^{2}$. Note that the sources are the direct 1D analogue of targets [3, 4] and topologically charged spirals do not exist in 1D.
@@
Figure 4: (a) Defect density $\rho_{D}$ versus $\alpha$ for P-BC shows nontrivial behaviour. It is low at certain values of $\alpha\simeq 0.08,0.2,0.3$, which we call “resonances” (shaded green). Overall it first increases with $\alpha$, peaks around $\alpha\sim 0.2$ and then starts to decrease. Above $\alpha_{c}\sim 0.6$, $\rho_{D}$ vanishes marking the onset of the travelling wave phase. $\alpha_{c}$ is close to the crossover threshold $\alpha_{\times}\sim 0.62$ predicted by the wavenumber selection. (b) Average polar order $\bar{J}$ versus $\alpha$ in the long time steady state. For small $\alpha$, $\bar{J}\sim 0$, at large $\alpha$, $\bar{J}\sim 1$. Near $\alpha_{c}$ (grey shaded region), $\bar{J}$ shows large fluctuations due to finite size effects. Inset: The fraction of simulations $(F_{T})$ that show defect states for $\alpha$ close to $\alpha_{c}$. Below $\alpha_{c}$, $F_{T}=1$, above $\alpha_{c}$, all simulations transition to travelling waves, thus $F_{T}=0$. (c) Time ($t_{\text{st}}$) after which the system reaches the steady-state value of $\rho_{D}$. At resonances, defect merger events continue for long times, thus $t_{\text{st}}$ is orders of magnitude larger. To verify that the resonances are not the effect of domain size, we also plot results for $L=256\pi$ (black points), which shows similar behaviour. (d) Minimum inter-defect separation $d_{\mathrm{min}}$ versus $\alpha$. As defects keep merging for long times, we find sharp jumps in $d_{\mathrm{min}}$ in the resonance regions, which we also observe in 2D [4]. All plots share the same legend keys, and the location of the three resonances that are shaded green are at the same values of $\alpha\simeq 0.08,0.2,0.3$ for panels (a), (c), and (d).

### v2

# Disorder to Order Transition in 1D non-reciprocal Cahn-Hilliard Model

@@ Abstract
We present the phenomenology of the one dimensional non-reciprocal Cahn Hilliard model for varying non-reciprocity $(\alpha)$ and different boundary conditions. At small $\alpha$, a perturbed uniform state evolves to a defect laden configuration that lacks global polar order. Defects are the sources and sinks of travelling waves. For a given $\alpha$, defects with a unique wave number that increases monotonically with $\alpha$ are selected. A critical threshold $\alpha_{c}$ marks the onset of a transition to states with finite global polar order. For periodic boundary conditions, above $\alpha_{c}$, the system shows travelling waves that are completely ordered. In contrast, travelling waves are incompatible with the Neumann and Dirichlet boundary conditions. Instead, for $\alpha\gtrsim\alpha_{c}$, we find fluctuating domains that show intermittent polar order and at large $\alpha$, the system partitions into two domains with opposite polar order.
@@ I Introduction
The non-reciprocal Cahn-Hilliard (NRCH) model describes phase separation of multi-component mixtures in the presence of non-reciprocal couplings [1, 2, 3, 4, 5, 6, 7, 8]. Systems with non-reciprocity are intrinsically out of equilibrium and show rich phenomenology [9, 10, 11, 12, 13, 14, 15, 16]. In the NRCH model, the spontaneous breaking of parity and time-reversal symmetries leads to the formation of travelling density bands [1, 2], suppressed coarsening dynamics [17, 1], localized states [18], chaotic steady states [8], true long-range polar order in two dimensions [7], and enhanced stability in multispecies mixtures [5]. Introduced phenomenologically, the NRCH model emerges as a universal amplitude equation for a conserved-Hopf instability in systems with two conservation laws [19] and can also be derived from a systematic coarse-graining of a microscopic model of phoretically active Janus colloids [20, 21]. Non-reciprocal interactions emerge in a variety of non-equilibrium systems with hydrodynamic interactions [22] and vision-cone interactions [23, 24, 25], active mixtures [12, 26, 27, 14, 28, 15, 29], mixtures of active and passive components [30, 31], and some mass-conserved reaction-diffusion systems [32, 33].
@@
Studies of the NRCH model in two dimensions (2D) have revealed a rich phenomenology due to the presence of defects, namely, spirals with unit magnitude topological charge and topologically neutral targets [3, 4]. These defects are the sources of the travelling waves, and thus the precursors of global polar order. For a given strength of non-reciprocal coupling $\alpha$, defects with a unique asymptotic wave number, $k_{\infty}\propto\sqrt{\alpha}$ are selected [3]. At a critical threshold of nonreciprocity $\alpha_{c}$; we find a disorder-to-order transition. For $\alpha<\alpha_{c}$, random disordered states evolve to quasi-stationary defect networks with no global polar order. On the other hand, for $\alpha>\alpha_{c}$ the system admit travelling waves without any defects. Studying the interactions between pairs of defects reveals that the stability of targets in this model fundamentally alters the nature of dynamics of the defects [4], as compared with previously studied nonequilibrium systems with nonconserved order parameter [34, 35, 36, 37]. Thus, the NRCH model provides a fertile ground to investigate the defect dynamics in non-reciprocal systems with conserved order parameter.
In this paper, we present an extensive study on the defects of the NRCH model in one dimension (1D). As compared to our previous investigations in 2D [3, 4], which have showcased the complexity of defect dynamics in the NRCH model, the 1D configuration provides the opportunity to investigate the role of the defects in a simpler setting, for a number of reasons: (i) Topologically charged defects, which would be the analogue of spirals in 2D, are ruled out. (ii) Sources of travelling waves in 1D are the analogue of targets in 2D. Similarly, sinks are the 1D analogue of domain walls. (iii) Waves emanated from the sources end at the sinks; thus sources and sinks are arranged in an alternating manner on the 1D domain. Furthermore, the defect motion is also limited to the 1D line. (iv) Since 1D allows only for sources and sinks, defect interactions are also simplified. At small separations, a source-and-sink pair annihilates. At larger distances, they form stable configurations. These reasons afford us a deeper and more pedagogical representation of the interplay between defect dynamics and the emergent properties of the NRCH model.
In addition to these effects, the 1D geometry allows us to explore the role of boundary conditions in the dynamics of the NRCH model. While the disorder-to-order transition at higher $\alpha$ leads to travelling waves for periodic boundary conditions, pure unidirectional travelling waves are incompatible with Neumann and Dirichlet boundary conditions, which are more relevant to experimental systems. How the transition proceeds in such cases and what kind of phases emerge from these transitions remain unknown. We investigate the defect dynamics and the effect of the boundary conditions on the disorder-order transition in the 1D NRCH model under different boundary conditions. Our main results are as follows: (i) We find a transition from disordered, defect laden states to polar ordered states. (ii) Below the transition threshold ($\alpha<\alpha_{c}$), sources and sinks with a particular wavenumber are selected. (iii) For periodic boundary conditions with $\alpha>\alpha_{c}$, we find travelling waves with perfect global polar order. (iv) For Neumann and Dirichlet boundary conditions, the phenomenology above $\alpha_{c}$ is different. For $\alpha$ slightly above $\alpha_{c}$, the system shows intermittent behaviour, with fluctuating domains of polar order. For larger $\alpha$, the system splits into subdomains with opposing polar order. (v) The transition point $\alpha_{c}\sim 0.60$ coincides with the crossover value $\alpha_{\times}\sim 0.62$ predicted by the Eckhaus instability of the plane waves. This is in a stark contrast with 2D, where the transition point $\alpha_{c}\sim 0.28$ is much smaller than the Eckhaus threshold $\alpha_{\times}\sim 0.58$.
@@
The rest of the paper is organized as follows. In Section II, we present our model and methods. In Section III, we present our results for periodic boundary conditions. We highlight the features of the source and sink solutions, wave number selection, defect density, polar order, and the disorder-to-order transition. In Section IV, we consider Dirichlet and Neumann boundary conditions and discuss how the boundary conditions alter the phenomenology. We conclude the paper with a discussion and future perspectives.
@@ II Model and Methods
We consider a minimal model of two conserved scalar fields $\phi_{1}(x,t)$ and $\phi_{2}(x,t)$ with non-reciprocal interactions on a one dimensional domain. The conserved dynamics of the complex scalar order parameter $\phi=\phi_{1}+i\phi_{2}$ is governed by the continuity equation $\partial_{t}\phi=\partial_{x}^{2}\mu$, where $\mu=\delta\mathcal{F}/\delta\phi^{*}+i\alpha\phi$ is the non-equilibrium chemical potential. The equilibrium contribution to $\mu$ promotes phase-separation and is derived from the free energy functional $\mathcal{F}=\int\mathop{}\!\mathrm{d}x\left(-|\phi|^{2}/2+|\phi|^{4}/4+|\nabla\phi|^{2}/2\right)$. The nonequilibrium part encodes the non-reciprocal interactions between the two species; quantified by the parameter $\alpha$. For $\alpha=0$, the model describes equilibrium phase-separation of two interacting species, and an initial disordered state coarsens towards a bulk-phase separated state. Any $\alpha\neq 0$ sets the system out of equilibrium and it describes non-reciprocal phase-separation. A positive $\alpha$ implies that $\phi_{1}$ chases $\phi_{2}$, and the model is invariant under the simultaneous transformation $\alpha\to-\alpha$ and $\phi_{1,2}\to\phi_{2,1}$. Combining these terms, we obtain the following non-dimensional equation for the evolution of $\phi(x,t)$ [3, 4, 1, 2]
@@
which reveals that $\forall~|k|<1$ the perturbations grow in an oscillatory manner. The behaviour of the system depends on $\alpha$, domain length $L$, and the choice of boundary conditions. We systematically study the phase-behaviour of Eq. 1 with varying strength of nonreciprocity $(\alpha)$ for various domain sizes $(L)$ and boundary conditions. We consider three different kinds of boundary conditions that are relevant to the phenomenon of phase separation [38]:
@@
    Neumann boundary conditions (N-BC), where
    $\partial_{x}\phi|_{0}=\partial_{x}\phi|_{L}=0~\mathrm{and}~\partial_{x}\mu|_{0}=\partial_{x}\mu|_{L}=0.$ (4)
@@
    Dirichlet boundary conditions (D-BC), where we fix $\phi$ on boundaries and $\mu$ is flux-free as
    $\phi(0)=\phi(L)=0~\mathrm{and}~\partial_{x}\mu|_{0}=\partial_{x}\mu|_{L}=0.$ (5)
One can also consider fixed boundary conditions for $\mu$, but it is not compatible with the conservation of mass, thus we do not consider this case.
Figure 1: Kymograph plot of the polar order parameter $J(x,t)$ highlighting the evolution of disordered states. Defects are the zeros of $J(x,t)$, which appear white in the kymograph. We also mark the exact defect positions (small black points) at late times. Initially, numerous source-sink pairs are formed, which merge and the system settles into a stable configuration, which persists till the end of the simulations.
Figure 2: Sources and sinks in the 1D NRCH model. (a, b) Representative plot of $\real{\phi}$ and the polar order parameter $J$ (normalized by its maximum value), respectively. Sources (blue circles) and sinks (orange circles) are arranged in an alternating manner on the 1D domain. $J$ vanishes at defect centers and is constant far from defects, where the wavefront is that of a plane wave. (c, d) Plot of $R(x)$ and $k(x)$ for sources and sinks for different values of $\alpha$. For these plots, $x$ is measured from the center of the defects. At defect centers, $\partial_{x}R$ and $k(x)$ vanish. Far from the defects, $R(x)$ and $k(x)$ approach their plane wave forms.
To study the phenomenology, we numerically integrate Eq. 1 on a 1D domain of length $L$ that is discretized using $N$ points. For P-BC, we use a pseudo-spectral algorithm with a second-order exponential time-differencing scheme [39] for time marching. For N-BC and D-BC, we perform the simulations using a Chebyshev polynomials based spectral Tau method implemented by the Dedalus framework [40]. Additionally, we use Dedalus to cross-check our results for the periodic domain. The source code to simulate the 1D NRCH model for all the boundary conditions considered here using Dedalus is available at https://github.com/navdeeprana/DedalusScripts.
We start all our simulations with a small perturbation to the homogeneous state. i.e., $\phi(x,0)=\delta\phi(x)$, where at every point $x$ the initial perturbation $\delta\phi(x)$ is drawn from a uniform distribution. Further, we impose $\int_{0}^{L}\phi(x,0)\mathop{}\!\mathrm{d}x=0$, which remains constant throughout the entire time evolution due to the conservation of mass. Unless reported otherwise, we average over eight or more independent realizations of initial conditions for each data point reported in the paper.
@@ III Periodic boundary conditions
For P-BC, the NRCH model (1) admits travelling wave solutions of the form
@@
where $k<1$, $R^{2}=1-k^{2}$, and $\omega=\alpha k^{2}$. For small $k$, travelling states are stable to small perturbations, however an Eckhaus instability kicks in at large $k$ and restricts the allowed wave number range to $0\leq k^{2}<1/3$ [35, 8, 3, 41]. In our previous studies [3, 4], we have shown that in 2D, at low $\alpha$, disordered states evolves to defect configurations composed of spirals and targets. In 1D, we find similar phenomenology. A random state evolves into multi-defect configurations that persist up to the end of simulation time (see Fig. 1 and Fig. 2). We now discuss the properties of these states in detail.
Figure 4: (a) Defect density $\rho_{D}$ versus $\alpha$ for P-BC. Overall $\rho_{D}$ peaks around $\alpha\sim 0.2$ and vanishes for $\alpha\to 0$ and $\alpha\to\alpha_{c}$. Above $\alpha_{c}\sim 0.6$, $\rho_{D}=0$ which implies the travelling wave phase. At certain values of $\alpha\simeq 0.08,0.2,0.3$, which we call “resonances” (shaded green), $\rho_{D}$ significantly dips. (b) Average polar order $\bar{J}$ versus $\alpha$. For small $\alpha$, $\bar{J}\sim 0$ and at large $\alpha$, $\bar{J}\sim 1$. For $\alpha\sim 0$ or $\alpha\sim\alpha_{c}$ (grey shaded region), $\bar{J}$ shows large fluctuations due to finite size effects. We compute $\bar{J}$ separately for defect-riddled states (solid markers) and travelling states (hollow markers). In the shaded region, some realizations show defects, some show travelling waves. Inset: The fraction of ensemble $(F_{D})$ that show defect states for $\alpha$ close to $\alpha_{c}$. Below $\alpha_{c}$, $F_{D}=1$, above $\alpha_{c}$, all simulations transition to travelling waves, thus $F_{D}=0$. Larger domains show sharper transitions. From these plots, we infer $\alpha_{c}\sim 0.6$.
@@ III.1 Sources and Sinks
$\phi(x,t)=R(x)\,e^{i\left(Z(x)-\omega t\right)},$ (7)
@@
where $x$ is measured from defect location, $R(x)$ is the amplitude, $Z(x)$ is the phase, and $k(x)\equiv\mathop{}\!\mathrm{d}Z/\mathop{}\!\mathrm{d}x$. They are the sources and sinks of travelling waves and we shall collectively call them defects. In Fig. 2(a), we plot a representative multi-defect configuration obtained from our P-BC simulations. The sources and sinks are arranged in an alternating manner on the 1D domain and the number of sources is always equal to the number of sinks. It is easy to identify the defect locations from the polar order parameter
$J(x,t)=\mathrm{Im}(\bar{\phi}\partial_{x}\phi)=\frac{1}{2i}\left(\bar{\phi}\partial_{x}\phi-\phi\partial_{x}\bar{\phi}\right).$ (8)
@@
As shown in Fig. 2(b), at a source (sink), $J$ vanishes with a positive (negative) slope, thus at a given time $t$, the number of defects is equal to the number of zeros of $J(x,t)$. Quantities similar to $J(x,t)$ have been used earlier to analyse the defect dynamics in systems with complex order parameter [42]. Substituting the defect ansatz in the definition of $J(x,t)$, we further find that at the defect core $\partial_{x}(R^{2})=\partial_{x}(\phi\bar{\phi})=2\bar{\phi}\partial_{x}\phi=2\phi\partial_{x}\bar{\phi}$. Thus, at the defect positions,
$\partial_{x}R(x)\Big|_{x=0}=0,~\mathrm{and}~k(0)=0.$ (9)
@@
In Fig. 2(c,d), we plot the profiles of $R(x)$ and $k(x)$ for sources and sinks for different $\alpha$. For a source, $k(x)>0(<0)$ when $x>0(<0)$, and vice versa for a sink. It is evident that R(x)=R(x)R(-x)=R(x) and k(x)=k(x)k(-x)=-k(x), which implies that for small $x$, $R(x)\sim a_{0}-a_{2}x^{2}$ and $k(x)\sim b_{1}x-b_{3}x^{3}$. Far away from the defect center, the wavefront approaches that of a travelling wave meaning for $x\gg 1$, $k(x)\to\pm k_{\infty}$ for a source and $k(x)\to\mp k_{\infty}$ for a sink (see Fig. 2(a)). The amplitude $R(x)\to\sqrt{1-k_{\infty}^{2}}$ also becomes constant and to ensure proper oscillations, we require $\omega=\alpha k_{\infty}^{2}$. Note that the sources are the direct 1D analogue of targets [3, 4] and topologically charged defects do not exist in 1D.
@@
Figure 5: Analysis of resonances. (a) Time ($t_{\text{st}}$) after which the system reaches the steady-state value of $\rho_{D}$. At resonances, defect merger events continue for long times, thus $t_{\text{st}}$ is orders of magnitude larger. To verify that the resonances are not the effect of domain size, we also plot results for $L=256\pi$ (black points), which shows similar behaviour. (b) Minimum inter-defect separation $d_{\mathrm{min}}$ versus $\alpha$. As defects keep merging for long times, we find sharp jumps in $d_{\mathrm{min}}$ in the resonance regions.